Circular No. 7422

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1999 C2 and C/2000 J3 D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of two more Kreutz sungrazing comets found in SOHO LASCO C3 data. C/1999 C2 was found by Biesecker, and no tail is evident. C/2000 J3 was found by K. Černis and M. Boschat via the SOHO website; it is also visible in C2 data, where a tail is evident, and Černis estimates m_{1} ∼ 7 on May 9.9 UT. Astrometry (measured by Biesecker and D. Hammer, reduced by B. G. Marsden) and orbital elements appear on MPEC 2000-J43 and 2000-J44.

Comet  Date  UT  α_{2000}  δ_{2000}
C/1999 C2  1999 Feb. 6.099  21^{h}35.^{m}2  −17^{o}04′
C/2000 J3  2000 May 9.725  3 16.0  +15 33

SUPERNOVA 2000ce IN UGC 4195 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000ce (cf. IAUC 7417), taken by P. Berlind on May 10.14 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova after maximum light, exhibiting a relatively flat continuum with well-developed lines of Si II, S II, Fe II, Mg II, and Ca II. Adopting the NASA/IPAC Extragalactic Database recession velocity for the host galaxy (1591 km/s) yields an expansion velocity for the supernova of 7600 km/s, measured via the Si II (rest 635.5 nm) feature. The age of the supernova, based on its spectral features (Riess et al. 1997, A.J. 114, 722), is 6 ± 3 days after maximum light. Concurrent images of SN 2000ce taken by P. Groot at the FLWO 1.2-m telescope yield B ∼ 15.8 and V ∼ 15.1.

V4334 SAGITTARII R. Joyce and K. Hinkle, Kitt Peak National Observatory (KPNO),

report infrared photometry (uncertainty 0.1 mag or better) of V4334 Sgr (cf. IAUC 6322) obtained with narrowband filters at the KPNO 2.1-m telescope (+ Phoenix spectrograph in imaging mode) on Apr. 13.5 UT: [1.083 μm] = 13.8, [1.282 μm] = 13.6, [1.558 μm] = 11.0, [2.230 μm] = 7.0, [3.226 μm] = 4.0, [4.651 μm] = 2.2. V4334 Sgr is slightly fainter at K, and slightly brighter at L and M, than in 1999 Sept. (cf. IAUC 7266). The infrared photometry, except for [1.083 μm], is consistent with a 670-K blackbody.

2000 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green